Diatomic O2 can react with the element magnesium to form magnesium oxide (MgO). The balanced chemical equation is: O2 + 2 Mg → 2 MgO
If 4 moles of magnesium totally reacted with more than enough O2, how many moles of MgO would be expected to form?
1
Expert's answer
2020-05-04T01:52:10-0400

Solution.

"4mol" "x mol"

"O_2+2Mg\to 2MgO;"

"2" "2"

We make the proportion: "\dfrac{4mol}{2}=\dfrac{xmol}{2};"

"x=\dfrac{4mol\sdot2}{2}=4mol;"

"\nu(MgO)=4mol" ;

Answer: "\nu(MgO)=4mol."